THE INFRARED EMISSION SPECTRUM OF DCL

\underline{TREANA PAREKUNNEL}, TSUYOSHI HIRAO, PINA COLARUSSO, ROBERT J. LE ROY and PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1.

The vibrational rotational emission spectrum of DCl was recorded with a Fourier transform spectrometer; line positions for the 1-0, 2-1, 3-2 and 4-3 bands of D^{35}Cl and the 1-0, 2-1 and 3-2 bands of D^{37}Cl were measured. The two sets of data were combined with previously measured submillimeter wave pure rotational transitions to obtain Dunham constants. A combined isotopomer analysis of data for all four isotopic forms of HCl was also performed combining these new data with analogous literature results for H^{35}Cl and H^{37}Cl, to yield Born-Oppenheimer breakdown correction terms for this system. The results of similar analyses of HF and HBr allowed a comparison of Born-Oppenheimer breakdown correction terms for HF, HCl and HBr.